APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Saturday March 4, 2006 (MJD 53799) ---=== OBSERVING TEAM ===--- Elena Malanushenko with some help from H and V. (afternoon checkout) Viktor Malanushenko and Howard Brewingotn (night) Jon Brinkman (phone) Mark Klaene (phone) French Leger (phone) ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- At evening - strong clouds. But after 5:30Z was quite good sky! We are finished 6 plates and got signal on one other plate: 2128. There was some problems with DIMM etc., see in to 'log' & 'problems'. ---=== OBSERVING LOG ===--- Afternoon: ------------ We a couple of problems with the afternoon checkout. Because the tcc session at the laptop would not take track commands, we had to reboot the mcp. And the computer monitor, photometric 5, had a blank screen during the doDrift -bias. However, the monitor's operation was restored by turning the power switch off, waiting 30-seconds, and turning the power on again. Spectro focus was checked but not adjusted; see Problems Section for focus numbers. We can not check the dust value - sensor does'nt work, - will using 3.5m sensors. Night: ------------ 01:52Z Howard check dust counter manually: it is working, the value more than 2300! But still not connection with program. 02:13Z dust: >=2500 02:31Z Doors open, fans on, louvers open, The sky is overcast; we are waiting. 05:44Z Cart.1 Plate 2429 (chunk135); Guide stars on goToField. Dust cam work now! Plate was finished in three exposure, DONE. 07:02Z Cart.7 Plate 2016 (chunk100); Guide stars on goToField. Plate was finished in three exposure, DONE. 08:00Z Cart.4 Plate 2354 (chunk123); using FK5 star & clean offset. Plate was finished in three exposure, DONE. 09:07Z Cart.5 Plate 2222 (chunk119); Guide stars on goToField. Plate was finished on one exposure, accumulated S/N=19.1 (with one of previous night) - is DONE. 09:46Z Cart.6 Plate 2230 (chunk119); Guide stars on goToField. Plate was finished in three exposure, DONE. 10:42Z Cart.9 Plate 1703 (chunk 87); Guide stars on goToField. Plate was finished in three exposure, DONE. 11:40Z Cart.8 Plate 2128 (chunk109); Guide stars on goToField. Plate is NOT finished, we are got S/N=13.3 only. endNight finished and was happy ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- 6042 22:48Z 23:09Z 100 O -146.32 -141.01 49 ignore ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 37850, plate 2222 ------- 22:59 37850 10.0 flat 23:02 37851 2.0 arc 23:07 37852 2.1 arc 23:09 37853 2.1 arc ----- sequence 37854, plate 2429 ------- 23:09 37854 10.0 flat 05:49 37855 10.0 flat 05:52 37856 2.0 arc 06:10 37857 900.1 target 4.5 5.8 6.5 6.4 06:28 37858 900.1 target 4.8 6.2 7.2 7.4 06:50 37859 1200.1 target 7.8 9.0 11.5 9.3 ----- sequence 37860, plate 2016 ------- 07:05 37860 10.0 flat 07:07 37861 2.0 arc 07:26 37862 900.1 target 8.0 8.1 8.9 7.5 07:38 37863 600.1 target 5.0 5.3 5.1 4.7 07:51 37864 600.1 target 5.0 5.3 4.9 4.6 ----- sequence 37865, plate 2354 ------- 08:09 37865 10.0 flat 08:11 37866 2.0 arc 08:27 37867 720.1 target 4.6 5.6 5.1 5.9 08:41 37868 720.1 target 4.9 5.8 5.2 5.8 08:57 37869 780.1 target 5.3 6.2 5.9 6.3 ----- sequence 37870, plate 2222 ------- 09:10 37870 10.0 flat 09:12 37871 2.0 arc 09:36 37872 1200.1 target 11.9 10.2 12.9 11.2 ----- sequence 37873, plate 2230 ------- 09:48 37873 10.0 flat 09:51 37874 2.0 arc 10:06 37875 720.1 target 6.6 6.6 6.7 6.7 10:19 37876 600.1 target 5.2 5.1 5.0 5.4 10:31 37877 600.1 target 4.9 4.9 5.0 5.3 ----- sequence 37878, plate 1703 ------- 10:44 37878 10.0 flat 10:46 37879 2.0 arc 11:00 37880 600.1 target 4.6 4.7 3.8 4.9 11:17 37881 900.1 target 8.1 7.3 7.4 7.7 11:30 37882 600.1 target ----- sequence 37883, plate 2128 ------- 11:42 37883 10.0 flat 11:44 37884 2.0 arc 12:00 37885 720.1 target 5.8 4.9 5.3 4.7 12:15 37886 780.1 target 6.1 4.9 5.5 5.0 12:32 37887 822.8 target 3.7 4.0 3.8 3.6 ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ 05:58Z 1 2429 -42.19 0.0004 67.51 0.0028 144.11 -0.0086 1.000110 07:11Z 7 2016 105.26 0.0021 63.58 0.0058 -93.12 -0.0135 1.000230 08:14Z 4 2354 272.51 -0.0002 70.13 0.0013 104.50 -0.0054 1.000140 09:16Z 5 2222 275.27 0.0019 73.02 0.0021 105.37 -0.0165 1.000050 09:54Z 6 2230 272.83 -0.0013 70.97 0.0027 104.31 0.0022 1.000020 10:49Z 9 1703 -22.17 -0.0021 67.09 0.0032 161.01 0.0012 1.000240 11:59Z 8 2128 286.47 0.0009 68.11 0.0027 117.55 -0.0099 1.000130 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL Stays: JL ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 05:58Z 1 2429 1.00011 -1086 286 50 -42 67.5 6.2 14 265 ---- 1.2 07:11Z 7 2016 1.00023 -2271 -686 50 105 63.6 5.7 21 274 ---- 1.4 08:14Z 4 2354 1.00014 -1383 62 75 273 70.1 5.4 20 276 ---- 1.2 09:16Z 5 2222 1.00005 -494 818 100 275 73.0 5.1 16 270 ---- 1.2 09:54Z 6 2230 1.00002 -198 1066 100 273 71.0 4.7 19 270 ---- 1.2 10:49Z 9 1703 1.00024 -2370 -725 100 -22 67.1 4.5 12 280 ---- 1.2 11:59Z 8 2128 1.00013 -1284 172 100 286 68.1 3.9 13 283 ---- 1.2 ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:52Z 47 21 20 240 (WSW) -9999 - clouds! 23:22Z 47 20 18 263 (W) -9999 - clouds! 23:52Z 47 19 15 258 (WSW) -9999 - clouds! 00:47Z 46 18 13 281 (W) -9999 - clouds! 01:20Z 44 14 15 271 (W) -9999 - clouds! 01:52Z 44 11 10 270 (W) -9999 - 02:24Z 44 10 12 262 (W) -9999 - 02:56Z 43 10 10 282 (WNW) -9999 - 03:28Z 43 10 15 311 (NW) -9999 - 04:01Z 42 10 6 294 (WNW) -9999 - 04:33Z 42 9 16 293 (WNW) -9999 - 05:04Z 42 10 12 325 (NW) -9999 - 05:37Z 43 9 10 255 (WSW) 688 - 06:09Z 42 8 12 268 (W) 695 - 06:41Z 43 8 15 264 (W) 734 - 07:11Z 42 9 16 268 (W) 696 - 07:42Z 42 8 14 262 (W) 710 - 08:12Z 41 7 20 270 (W) 714 - 08:42Z 41 8 21 272 (W) 805 - 09:13Z 41 7 14 273 (W) 811 - 09:43Z 40 7 16 269 (W) 773 - 10:13Z 40 5 17 273 (W) 796 - 10:43Z 40 5 12 285 (WNW) 848 - 11:15Z 39 5 10 271 (W) 843 - 11:45Z 39 4 15 297 (WNW) 874 - 12:15Z 38 3 12 311 (NW) 798 - 12:45Z 38 3 14 291 (WNW) 852 - 13:16Z 38 2 15 302 (WNW) 881 - ---=== TELESCOPE STATUS ===--- 02:31Z Doors and louvers opened, fans on. 05:30Z Telescope open. 12:50Z Enclosure off 12:50Z Enclosure on, doors and louvers closed, fans off. Telescope at 121,30,00 mount; cart 8 mounted. LN2 connected and turned on; Spectro were force filled. Imager 165 lbs 21 psi Spectro 178 lbs 27 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_18_1 Watcher: v2_31_0 MCP: v5_26_2 TPM: tpm_v3_3_0 AstroDa: v15_11 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_11 SoS: v5_1_0 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.8520 -5.8440 Axial B -5.6880 -5.6770 Axial C 0.9400 0.9420 Trans D -10.0050 -9.9940 Lateral E 10.3759 10.3759 Lateral F 11.9507 11.9380 GALILS Commanded: 4700. -2450. 3000. -7800. 3300. 3000. Actual: 4711. -2457. 2997. -7825. 3296. 3024. SETMIR VALUES PriDesOrient: 0.00 -1.68 7.05 385.25 92.02 PriOrient: 0.00 -1.52 7.02 384.99 93.81 SECONDARY: ---------- Focus: 0.00 Air Temp.: 9.1 Alt.: 29.999981 MIGS TONIGHT NOMINAL Axial A 1.5410 1.5420 Axial B 1.0570 1.0470 Axial C 1.1710 1.1650 Trans D -0.1260 -0.1240 GALILS Commanded: 1593406. 1603214. 1554144. -6800. -5900. Actual: 1593447. 1603642. 1554316. -6820. -5882. SETMIR VALUES SecDesOrient: 1257.00 15.87 0.00 -50.00 136.76 SecOrient: 1256.83 15.82 -0.11 -49.63 136.77 ---=== PROBLEMS IN DETAIL ===--- ---=== Spectro Focus Numbers ===--- IDL 5.5a> collimate, 37852 % Compiled module: COLLIMATE. Reading /data/spectro/53799/sdR-b1-00037852.fit speclog_version: Command not found. Reading /data/spectro/53799/sdR-b1-00037853.fit speclog_version: Command not found. COLLIMATE: Min offset = -0.43 pix COLLIMATE: Max offset = 1.04 pix COLLIMATE: RMS across CCD = 0.42 pix COLLIMATE: Mean offset = 0.36 pix COLLIMATE: Camera b1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict blue camera ring movement of -13.0 degrees for spectro-1 COLLIMATE: (if red is already in focus) Log file = Collimate-53799-b1-00037852.log PostScript file = Collimate-53799-b1-00037852.ps Reading /data/spectro/53799/sdR-b2-00037852.fit speclog_version: Command not found. Reading /data/spectro/53799/sdR-b2-00037853.fit speclog_version: Command not found. COLLIMATE: Min offset = -0.07 pix COLLIMATE: Max offset = 0.74 pix COLLIMATE: RMS across CCD = 0.23 pix COLLIMATE: Mean offset = 0.30 pix COLLIMATE: Camera b2 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict blue camera ring movement of -10.8 degrees for spectro-2 COLLIMATE: (if red is already in focus) Log file = Collimate-53799-b2-00037852.log PostScript file = Collimate-53799-b2-00037852.ps Reading /data/spectro/53799/sdR-r1-00037852.fit speclog_version: Command not found. Reading /data/spectro/53799/sdR-r1-00037853.fit speclog_version: Command not found. COLLIMATE: Min offset = -0.69 pix COLLIMATE: Max offset = 0.19 pix COLLIMATE: RMS across CCD = 0.26 pix COLLIMATE: Mean offset = -0.29 pix COLLIMATE: Camera r1 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of 2617 steps for spectro-1 COLLIMATE: Issue SOP command: "sp1; mechPiston 2617" Log file = Collimate-53799-r1-00037852.log PostScript file = Collimate-53799-r1-00037852.ps Reading /data/spectro/53799/sdR-r2-00037852.fit speclog_version: Command not found. Reading /data/spectro/53799/sdR-r2-00037853.fit speclog_version: Command not found. COLLIMATE: Min offset = -0.77 pix COLLIMATE: Max offset = -0.08 pix COLLIMATE: RMS across CCD = 0.19 pix COLLIMATE: Mean offset = -0.32 pix COLLIMATE: Camera r2 appears to be OUT-OF-FOCUS (|mean| > 0.15 pix) COLLIMATE: Predict (red) piston movement of 2903 steps for spectro-2 COLLIMATE: Issue SOP command: "sp2; mechPiston 2903" Log file = Collimate-53799-r2-00037852.log PostScript file = Collimate-53799-r2-00037852.ps % Program caused arithmetic error: Floating illegal operand IDL 5.5a> ------------------------------- We lost a few minutes because we had to do the cals for plate 2429 twice; see below. Flat Field Median 666.8 Starting Flat field 0 (press cntrl-C to abort) Flushing charge for 15 sec appending BEGIN Exposure is paused; 10s of data were taken appending END EXPOSURE TIME is: 10 2006-03-05 05:46:00Z: NaN value for IPARATE in spectro fits header 1141537528 1141537557 Seq 37854 Exp 37854 Frame 37854 Do not schedule startFlushCharge sp1 200 Do not schedule startFlushCharge sp2 200 Finished with Flat field 0 Finished with Flat Field Closing Flat Field Screen 4 Ne and 4 HgCd lamps are on Starting Arc lamp exposure 0 (press cntrl-C to abort) goStare was not successful, trying again goStare was not successful, trying again goStare was not successful, trying again goStare was not successful, trying again goStare was not successful, trying again goStare was not successful, trying again Could not do goStare in 5 attempts Finished with Arc Lamp exposure 0 Finished With Arc Lamps Opened Flat Field Screen Exiting successfully This calibration took 153.672 seconds sp1> Also in SOS, we got the following warning. sdR-b1-00037854.fit: ABORT: Unable to reduce flat ------------------- We had a problem with the DIMM tonight. Although the software was taking images of Polaris, it would not output any seeing data. I did not have time to do any trouble-shooting but will perform some checks on Sunday evening. Later in the evening, though, the DIMM started displaying the seeing data. Yet, the data did not look normal. We will check the DIMM more closely on Sunday evening. ------------------- At evening does not worked Dust Camera, it start working after 05:30Z only...